# Barriers that Programming Instructors Face While Performing Emergency Pedagogical Design to Shape Student-AI Interactions with Generative AI Tools

### v1

@@ 3.4. Survey
Since our interviewees were selected specifically because they had already conducted interventions to improve student-AI interactions, we recognized that our interviews would likely emphasize perspectives of instructors who wanted to keep including GenAI usage as part of their courses. In order to reduce the impact of this limitation, in parallel with our interviews we conducted a survey to gather perspectives of computing faculty more broadly, including faculty that had not changed their course materials in response to GenAI usage. The survey was designed to be completed within 10 minutes and asked participants questions about their personal perspectives and perspectives of their colleagues on GenAI usage in their courses. We also included demographic questions, such as average number of courses taught per year and number of students taught per year. The survey was distributed via email lists for computing faculty, discussion forums, and personal outreach. To capture viewpoints from more diverse faculty, we specifically reached out to email lists for faculty at Minority-Serving Institutions (MSIs) (liCharacteristics2007, 37) and Historically Black Colleges and Universities (HBCUs) (allenHistorically2007, 3). A total of 169 faculty responded to the survey. More than half of these faculty taught at MSIs (51%), and a sizable proportion taught at HBCUs (17%). Most faculty either taught 1-2 courses per academic term (37%) or 3-4 courses per academic term (41%), although a small minority taught more than 5 courses per term (17%).
@@ 3.5. Data Analysis
During each interview, the researcher who conducted the interview took timestamped notes. Then, a second researcher independently listened to each Zoom recording and took their own set of notes. Both researchers met weekly to discuss their notes together. After the interviews were completed, we iteratively came up with themes using an inductive approach (corbinBasics2014, 9) – we annotated interview quotes with potential themes, then merged and split themes based on our weekly discussions until we converged on our final set. During this process, we decided to center each of our themes on a barrier that instructors face while attempting to alter student-AI interactions because of the prevalence of repeated barriers across our interviewees.
To analyze the survey data, two researchers computed counts and proportions for Likert-scale questions and applied an inductive approach to generate themes for open-ended responses. Although we allowed survey participants to leave any question blank if they chose, we found that all of our Likert-scale questions had response rates over 90% (between 152-161 responses). As such, when reporting Likert-scale question results, we omit the number of responses for ease of reading.
@@ 3.6. Study Scope and Limitations
The survey used convenience and snowball distribution via email lists, forums, and personal outreach rather than probability sampling. We did not weight responses and allowed missing items, so we cannot estimate population parameters or make inferential claims beyond respondents. Although we sought to include a more diverse range of responses from MSIs and HBCUs, the sample is nonetheless not nationally representative. Because of these limitations, in this paper we treat survey results as descriptive of our respondents and use percentages mainly to compare or contrast with interview patterns rather than as prevalence estimates.
@@ 4. Results
Table 2. Summary of the five barriers in engaging in emergency pedagogical design, surfaced through our interviews with computing instructors.
@@
In this section we present a definition of emergency pedagogical design in the context of instructors adapting their courses to account for student GenAI usage (Section 4.1). During this design work, instructors faced five barriers that repeatedly arose during our analysis: fragmented buy-in (Section 4.2); policy crosswinds (Section 4.3); implementation challenges (Section 4.4); assessment misfit (Section 4.5); and lack of resources (Section 4.6).
@@ 4.1. Overview of Emergency Pedagogical Design
In the context of our instructors reacting to student usage of GenAI tools, we define emergency pedagogical design as the work that instructors do to adjust their courses in order to improve student interactions with GenAI. In our study, we observed the following properties of emergency pedagogical design:
@@
  • Indirect rather than direct: Interaction designers typically have the ability to make direct changes to the interface they are designing. In contrast, instructors lack the ability to directly update interfaces of GenAI tools, since these tools are maintained by corporations rather than academic institutions. In addition, GenAI output is increasingly embedded within everyday interfaces (e.g., web search (reidGenerative2024, 54) and code editors (hollandAnnouncing2024, 27)), making it challenging for instructors to anticipate all the ways that students can trigger GenAI usage, whether intentionally or inadvertently. Without the ability to change GenAI interfaces directly, instructors indirectly influence student interactions with GenAI through learning materials, course policies, assessment strategies, and course infrastructure.
@@
Taken together, these properties portray emergency pedagogical design as reactive, indirect, and iteration-driven work conducted under uncertainty and with limited visibility into student behavior. In the following subsections, we elaborate on five recurring barriers that shape and often constrain how instructors carry out this work, summarized in Table 2.
@@ 4.2. Fragmented Buy-In
We use buy-in to mean two levels of commitment: an instructor’s own motivation to revise course materials to shape student-AI interactions, and departmental support from colleagues that enables and rewards those revisions. Emergency pedagogical design competed with routine teaching work. Instructors were already committed to preparing and updating assignments, delivering lecture, meeting with staff, answering email, writing exams, and grading. Managing unprompted student use of generative AI arrived on top of these duties, not in place of them. As a result, designing student-AI interactions was rarely their top priority, even when instructors believed it mattered for learning.
Although we selected for interviewees that had already made concrete changes to their course materials in response to student GenAI usage, several instructors pointed out an initial lack of personal buy-in when they first learned about GenAI (P03-P06, P09-P10, P12). For example, one instructor mentioned that “I’d been trying to ignore [GenAI tools] for as long as I possibly could” by using proctored quizzes to deter students from using GenAI on assignments (P06). Despite having course policies forbidding GenAI, this instructor eventually realized that students used GenAI to complete assignments anyway, and then did poorly on proctored assessments – a pattern echoed by other interviewees (P04, P08, P12, P13). On reflection, P06 shared that “punishing them [via a proctored assessment] after the fact is not particularly useful or effective. So how can I get them to stop doing things they shouldn’t be doing?” Other instructors cited a desire to curb potential “bad habits” when students used GenAI without instructor guidance (P07, P08).
Beyond personal buy-in, departmental buy-in was also fragmented. Interviewees reported that they were in a small minority within their departments who were meaningfully integrating AI into coursework (P01, P05, P07, P08, P11), and sometimes even faced opposition from other faculty in their department about trying to include AI interactions in coursework (P05, P10-P12). The mixed reactions by other faculty in their departments reduced informal support, slowed coordination across multi-section courses, and made it harder to share materials or run aligned evaluations. This isolation also constrained scope: all of our interviewees piloted course-specific changes, but were not yet able to attain broader alignment with other courses via shared policies, common assignment patterns, or cross-course scaffolds.
@@
0%2525%5050%7575%100100% (a) How open are you to adopting GenAI technologies in your curriculum? (b) How open do you think your colleagues are to adopting GenAI in their curriculum? (c) My colleague’s resistance to GenAI is a barrier to wider adoption within the department. Not at all openNot at all openStrongly disagreeVery openVery openStrongly agree38204989626773111234250268
@@
Figure 2. Although a strong majority of survey respondents considered themselves open to adopting GenAI technologies (a), only a minority felt their colleagues were equally open (b). However, only a few instructors perceived colleague resistance to GenAI as a barrier (c), suggesting that ambivalence was more common than direct opposition. Instructors responded on a 5-point scale. The number of responses for each option is displayed within each bar.:
This fragmentation was echoed outside of our interviewees – although 81% of our survey respondents said that they were either Open or Very Open to adopting AI tools in their teaching, only 28% said the same about their colleagues  (Figure 2a,b). Instructors reported colleague ambivalence rather than direct opposition, since only 23% Agreed or Strongly Agreed with the statement: “My colleagues’ resistance to using GenAI tools in their courses is a barrier to wider adoption at my institution.” (Figure 2c).
@@ 4.3. Policy Crosswinds
Nearly all interviewees reported that neither their department nor institution provided prescriptive policies governing GenAI use in teaching as a result of fragmented buy-in (P01-P08, P10-P13). In the absence of top-down policies, instructors had to decide their own policies on a per-course basis. As a result, practices diverged, sometimes dramatically. As one instructor summarized: “From a student perspective, it’s the wild west. Some courses allow GenAI usage, some don’t.” (P07). Instructors pointed out that this caused “lots of confusion” (P01) for students who had to navigate different policies for every course that they took. Although several instructors mentioned departmental meetings to discuss GenAI policies, discussions had not yet converged shared guidelines or consistent approaches (P01, P02, P04, P05, P07, P10, P12). These inconsistencies also resulted in equity implications that current policies often did not address. For example, no instructors reported course policies that distinguished between access to paid versus unpaid tools or between standalone chatbots and GenAI embedded in everyday software (e.g., code editor assistants or web search).
As one notable counterexample, P09 described a university “two-lane” policy that distinguished proctored from unproctored assessments. For proctored assessments, instructors could set GenAI rules; for unproctored assessments like take-home programming exercises, “students can use whatever tools they choose, and instructors cannot prohibit GenAI use” (P09). This instructor voiced support for this approach as a pragmatic way to reduce rule conflict at the course level while aligning official policy with what could be reasonably enforced outside proctored settings.
The survey echoed these crosswinds. Knowing that students may use GenAI regardless of policies, 85% of respondents said they were Likely or Very Likely to change their curriculum (Figure 3a). By contrast, when asked how institutional constraints or policies affected their ability to integrate GenAI, only 26% reported a Positive or Very Positive impact (Figure 3b).
@@
0%2525%5050%7575%100100% (a) Knowing that students may use GenAI regardless of policies, how likely are you to make changes to your curriculum? (b) How do institutional policies impact your ability to integrate GenAI into the curriculum? Very unlikelyVery negativelyVery likelyVery positively51759752832562812
@@
Figure 3. Surveyed instructors felt motivation to update their curriculum in response to student usage of GenAI (a). Institutional constraints had mixed perceived impact (b). Instructors responded on a 5-point scale. The number of responses for each option is displayed within each bar. :
@@ 4.4. Implementation Challenges
0%2525%5050%7575%100100% (a) How important do you think it is to integrate GenAI tools into your curriculum? (b) How frequently do you use GenAI tools in course activities? Not at all importantNeverVery importantVery often352552762732402930
Figure 4. Surveyed instructors felt it was important to integrate GenAI tools into their curriculum (a), but fewer instructors reported actually doing so (b). Instructors responded on a 5-point scale. The number of responses for each option is displayed within each bar.:
@@ 4.6. Lack of Resources
Our interviewees dedicated substantial amounts of time, staff hours, and even money in the process of emergency pedagogical design. For example, bespoke infrastructure required ongoing development and maintenance. As mentioned in Section 4.4, P04 deployed a custom chatbot and P09 built an AI-graded explanation workflow; both required design, implementation, bug fixes, and upkeep during the term. These two instructors reported that these systems were only possible with support from software engineers and PhD students.
@@
Alternative assessments often meant a shift away from autograded work towards qualitative work, which also required more course staff to manage. As mentioned in Section 4.5, P02 reorganized his course around weekly stand-up meetings with every student. To handle this for his course of 300 students, he needed “a lot of TAs!” (P02). In practice, this course needed to recruit and train over 50 course staff members to cover approximately 300 students, a 6:1 staff to student ratio. Others emphasized the hand-grading burden for creative work, especially in courses that traditionally relied on autograder tests (P05, P08, P14).
@@
What is notable about our interviewees is that they taught at most two courses per academic term, and many had advantages such as partnerships, external funds, curriculum materials, or the ability to hire many course staff members. This pattern is striking: substantial changes to student-AI interaction were made by instructors with lighter teaching loads and access to staffing or funding.
In contrast to our interviewees, a higher proportion of our survey respondents had heavier teaching loads: 61% taught three or more courses per academic term. Among instructors at minority-serving institutions, this share rose to 70%. In the most extreme cases, 10 respondents reported teaching six or more courses per academic term; all were from MSIs. As one might expect, a similar proportion expressed a lack of resources: when asked if they had time to learn and integrate GenAI given their current workload, 62% of surveyed instructors responded Strongly Disagree, Disagree, or Neither (Figure 5a). When asked how well their institution supports faculty in adopting emerging technologies like GenAI, a nearly identical 63% of faculty responded Very Poorly, Poorly, or Neither (Figure 5b). These results underscore both resource requirements and constraints that can prevent instructors from making fruitful attempts to improve student-AI interactions.
0%2525%5050%7575%100100% (a) Do you feel you have the time to learn and integrate GenAI into your teaching, given your current workload? (b) How well does your institution support faculty in adopting GenAI? Strongly disagreeVery poorlyStrongly agreeExcellent15345035251327613623
Figure 5. A substantial number of our survey respondents expressed that their current workloads restricted their time to learn and integrate GenAI (a), and that their institutions did not support this work (b). Instructors responded on a 5-point scale. The number of responses for each option is displayed within each bar.:
Key Insights: Across interviews, the binding constraints were clear and separable: time to design and maintain changes, staffing to run labor-intensive assessments, and money for model access. Instructors who redesigned assessments or built tooling tended to have lighter teaching loads and additional support. Survey results align with this picture: most faculty reported heavier teaching loads (especially at MSIs) and also less capacity to integrate GenAI into their course materials.

### v2

@@ 3.4. Survey
Since our interviewees were selected specifically because they had already conducted interventions to improve student-AI interactions, we recognized that our interviews would likely emphasize perspectives of instructors who wanted to keep including GenAI usage as part of their courses. In order to reduce the impact of this limitation, in parallel with our interviews we conducted a survey to gather perspectives of computing faculty more broadly, including faculty that had not changed their course materials in response to GenAI usage. The survey was designed to be completed within 10 minutes and asked participants questions about their personal views, their perceptions of colleagues, and basic teaching demographics such as average number of courses taught per year and number of students taught per year.
The survey was distributed via email lists for computing faculty, discussion forums, and personal outreach. To capture viewpoints from more diverse faculty, we specifically contacted email lists for faculty at Minority-Serving Institutions (MSIs) (Li and Carroll, 2007) and Historically Black Colleges and Universities (HBCUs) (Allen et al., 2007). A total of 169 faculty responded to the survey. More than half of these faculty taught at MSIs (51%), and a sizable proportion taught at HBCUs (17%). Most faculty either taught 1-2 courses per academic term (37%) or 3-4 courses per academic term (41%), although a small minority taught more than 5 courses per term (17%). Because the survey relied on convenience and snowball sampling across lists with unknown membership sizes, and because respondents could skip any item, we could not calculate a response rate; instead, we treat the results as descriptive of respondents rather than as estimates of wider faculty attitudes.
@@ 3.5. Data Analysis
During each interview, the researcher who conducted the interview took timestamped notes. Then, a second researcher independently listened to each audio recording and took their own set of notes. Both researchers met weekly to discuss their notes together. After the interviews were completed, we iteratively came up with themes using an inductive approach (Corbin and Strauss, 2014). Both researchers independently annotated interview excerpts at the level of short utterances or question-answer turns, depending on where a shift in meaning occurred. We then compared annotations in weekly meetings. Coding proceeded as a negotiated process rather than through formal inter-rater reliability statistics: disagreements were discussed case by case, and codes were revised until both researchers reached consensus. This process unfolded over five weekly iterations, during which we merged, split, and renamed codes until the codebook stabilized. After convergence, we re-reviewed earlier excerpts to ensure consistent application. This iterative process led us to center each theme on barriers instructors encountered while attempting to alter student-AI interactions.
To analyze the survey data, two researchers computed counts and proportions for Likert-scale questions and applied an inductive approach to generate themes for open-ended responses. For open-ended items, both researchers independently labeled sentences within each response. We compared labels in weekly meetings and resolved disagreements through discussion until agreement was reached. Coding stabilized after three iterations, after which we re-reviewed earlier responses for consistency and grouped the agreed-upon codes into the themes reported in the results. Although we allowed survey participants to leave any question blank if they chose, we found that all of our Likert-scale questions had response rates over 90% (between 152-161 responses out of 169 submissions). As such, when reporting Likert-scale question results, we omit the number of responses for ease of reading.
@@ 3.6. Study Scope and Limitations
The survey used convenience and snowball distribution via email lists, forums, and personal outreach rather than probability sampling. We did not weight responses and allowed missing items, so we cannot estimate population parameters or make inferential claims beyond respondents. This approach likely introduced self-selection biases: instructors with strong views about GenAI, whether positive or negative, may have been more motivated to respond. Although we sought to include a more diverse range of responses from MSIs and HBCUs, the sample is nonetheless not nationally representative. Because of these limitations, in this paper we treat survey results as descriptive of our respondents and use percentages mainly to compare or contrast with interview patterns rather than as prevalence estimates.
Because all interviewees taught programming-focused courses and the survey focused on computing faculty, our findings are more directly relevant within computing education and may not generalize to other domains. We attempted to frame our findings so that they have the potential to appear in other disciplines, but we acknowledge that instructors in writing-intensive, humanities, or social science contexts may face distinct challenges that our data do not capture. As such, transferability of these findings beyond computing requires caution, and we view this work as an early case in one discipline where GenAI impacted teaching practices rapidly.
@@ 4. Results
Table 2. Overview of the five barriers that emerged from our analysis of interview and survey data. The barriers summarize recurring constraints computing instructors encountered while engaging in emergency pedagogical design.
@@
In this section we present a definition and properties of emergency pedagogical design as instructors adapted their courses to account for student GenAI usage (Section 4.1). During this design work, instructors faced five barriers that repeatedly arose during our analysis: fragmented buy-in (Section 4.2); policy crosswinds (Section 4.3); implementation challenges (Section 4.4); assessment misfit (Section 4.5); and lack of resources (Section 4.6).
Both the properties of emergency pedagogical design and the five barriers reported in this section were derived from the same inductive thematic analysis described in Section 3. During coding, we noted two kinds of recurring patterns: cross-cutting characteristics of the design setting that we refer to as properties of emergency pedagogical design, and recurring challenges that we grouped into five barriers. The properties characterize the conditions under which instructors worked, while the barriers describe the constraints that shaped their attempts to change student-AI interactions.
@@ 4.1. Overview of Emergency Pedagogical Design
In the context of our instructors reacting to student usage of GenAI tools, we define emergency pedagogical design as the work that instructors do to adjust their courses in order to improve student interactions with GenAI. In synthesizing interview accounts, we observed several cross-cutting patterns that described the broader conditions under which instructors tried to adjust their courses. We refer to these as properties of emergency pedagogical design because they characterize the design setting that emerged from our inductive analysis rather than specific challenges or solutions. We observed the following properties:
@@
  • Indirect rather than direct: Interaction designers typically have the ability to make direct changes to the interface they are designing. In contrast, instructors lack the ability to directly update interfaces of GenAI tools, since these tools are maintained by corporations rather than academic institutions. In addition, GenAI output is increasingly embedded within everyday interfaces (e.g., web search (Reid, 2024) and code editors (Holland, 2024)), making it challenging for instructors to anticipate all the ways that students can trigger GenAI usage, whether intentionally or inadvertently. Without the ability to change GenAI interfaces directly, instructors indirectly influence student interactions with GenAI through learning materials, course policies, assessment strategies, and course infrastructure.
@@
Taken together, these properties portray emergency pedagogical design as reactive, indirect, and iteration-driven work conducted under uncertainty and with limited visibility into student behavior. In the following subsections, we elaborate on five recurring barriers that constrained how instructors carried out emergency pedagogical design within this setting, summarized in Table 2.
@@ 4.2. Fragmented Buy-In
0%2525%5050%7575%100100% (a) How open are you to adopting GenAI technologies in your curriculum? (b) How open do you think your colleagues are to adopting GenAI in their curriculum? (c) My colleague’s resistance to GenAI is a barrier to wider adoption within the department. Not at all openNot at all openStrongly disagreeVery openVery openStrongly agree8204989626773111234250268
@@
Figure 1. Although a strong majority of survey respondents considered themselves open to adopting GenAI technologies (a), only a minority felt their colleagues were equally open (b). However, only a few instructors perceived colleague resistance to GenAI as a barrier (c), suggesting that ambivalence was more common than direct opposition. Instructors responded on a 5-point scale. The number of responses for each option is displayed within each bar.
Stacked bar plots showing responses to three survey questions: (a) How open are you to adopting GenAI technologies in your curriculum?; (b) How open do you think your colleagues are to adopting GenAI in their curriculum?; and (c) My colleague’s resistance to GenAI is a barrier to wider adoption within the department.
We use buy-in to refer to two levels of commitment: an instructor’s own motivation to revise course materials to shape student-AI interactions, and departmental support from colleagues that enables and rewards those revisions. We make this distinction because our survey results indicated a clear gap between instructors’ own openness to adopting GenAI and their perceptions of departmental support. While 81% of respondents rated themselves as Open or Very Open to adopting GenAI technologies, only 28% said the same of their colleagues (Fig. 1a,b). Although 23% Agreed or Strongly Agreed that colleague resistance was a barrier, most respondents selected neutral options, suggesting ambivalence rather than direct opposition (Fig. 1c). These patterns indicate that instructors often felt individually willing to revise courses but did not perceive a shared departmental commitment to this work.
Since we selected for interviewees that had already made concrete changes to their course materials in response to student GenAI usage, a limitation of our interviews is that they only surface perspectives of instructors who had strong personal buy-in for making course changes. Even so, several interviewees described an initial period of reluctance or skepticism (P03-P06, P09-P10, P12). For example, one instructor mentioned that “I’d been trying to ignore [GenAI tools] for as long as I possibly could” by using proctored quizzes to deter students from using GenAI on assignments (P06). Despite having course policies forbidding GenAI, this instructor eventually realized that students used GenAI to complete assignments anyway, and then did poorly on proctored assessments – a pattern echoed by other interviewees (P04, P08, P12, P13). On reflection, P06 shared that “punishing them [via a proctored assessment] after the fact is not particularly useful or effective. So how can I get them to stop doing things they shouldn’t be doing?” Other instructors cited a desire to curb potential “bad habits” when students used GenAI without instructor guidance (P07, P08).
Beyond personal buy-in, departmental buy-in was also fragmented. Interviewees reported that they were in a small minority within their departments who were meaningfully integrating AI into coursework (P01, P05, P07, P08, P11), a pattern that aligned with our survey data. Some even faced opposition from other faculty in their department about trying to include AI interactions in coursework (P05, P10-P12). The mixed reactions by other faculty in their departments reduced informal support, slowed coordination across multi-section courses, and made it harder to share materials or run aligned evaluations. This isolation also constrained scope: all of our interviewees piloted course-specific changes, but were not yet able to attain broader alignment with other courses via shared policies, common assignment patterns, or cross-course scaffolds.
@@ 4.3. Policy Crosswinds
0%2525%5050%7575%100100% (a) Knowing that students may use GenAI regardless of policies, how likely are you to make changes to your curriculum? (b) How do institutional policies impact your ability to integrate GenAI into the curriculum? (c) Unequal access to paid generative AI tools among students could create or exacerbate disparities in learning outcomes & workforce preparedness. Very unlikelyVery likelyVery negativelyVery positivelyStrongly disagreeStrongly agree5175975283256281256244680
@@
Figure 2. Surveyed instructors felt motivation to update their curriculum in response to student usage of GenAI (a). Institutional constraints had mixed perceived impact (b), yet unequal access to paid GenAI tools was a widely held concern (c). Instructors responded on a 5-point scale. The number of responses for each option is displayed within each bar.
Stacked bar plots showing responses to three survey questions: (a) Knowing that students may use GenAI regardless of policies, how likely are you to make changes to your curriculum?, (b) How do institutional policies impact your ability to integrate GenAI into the curriculum?, (c) To what extent do you agree or disagree with the following statement: ”Unequal access to paid generative AI tools among students could create or exacerbate disparities in learning outcomes and workforce preparedness.”
Since a growing number of institutions have started to issue guidelines about GenAI usage in coursework (McDonald et al., 2025), our survey asked computing instructors about the effects of these top-down policies on their adoption of GenAI. Knowing that students may use GenAI tools regardless of policies forbidding them, 85% of respondents said they were Likely or Very Likely to change their curriculum (Figure 2a). By contrast, when asked how institutional constraints or policies affected their ability to integrate GenAI, only 26% reported a Positive or Very Positive impact (Figure 2b). Respondents also expressed concerns about the uneven enforceability of institution-level policies, noting in open-ended responses that policies rarely specified how GenAI should be incorporated into assignments, how to interpret allowed versus disallowed uses, or how students’ access to paid tools should be handled.
Likewise, nearly all interviewees reported that neither their department nor institution provided prescriptive policies governing GenAI use in teaching as a result of fragmented buy-in (P01-P08, P10-P13). In the absence of top-down policies, instructors had to decide their own policies on a per-course basis. As a result, practices diverged, sometimes dramatically. As one instructor summarized: “From a student perspective, it’s the wild west. Some courses allow GenAI usage, some don’t.” (P07). Instructors pointed out that this caused “lots of confusion” (P01) for students who had to navigate different policies for every course that they took. Although several instructors mentioned departmental meetings to discuss GenAI policies, discussions had not yet converged shared guidelines or consistent approaches (P01, P02, P04, P05, P07, P10, P12). These inconsistencies also resulted in equity implications that current policies often did not address. For example, none of our interviewees reported course policies that distinguished between access to paid versus unpaid tools or between standalone chatbots and GenAI embedded in everyday software (e.g., code editor assistants or web search). The difference between paid and unpaid tools in particular was salient among survey respondents: 78% of responses Agreed or Strongly Agreed with the statement, “Unequal access to paid generative AI tools among students could create or exacerbate disparities in learning outcomes and workforce preparedness” (Figure 2c).
As one notable counterexample, P09 described a university “two-lane” policy that distinguished proctored from unproctored assessments. For proctored assessments, instructors could set GenAI rules; for unproctored assessments like take-home programming exercises, “students can use whatever tools they choose, and instructors cannot prohibit GenAI use” (P09). This instructor voiced support for this approach as a pragmatic way to align policy with what could be reasonably enforced outside proctored settings.
@@ 4.4. Implementation Challenges
0%2525%5050%7575%100100% (a) How important do you think it is to integrate GenAI tools into your curriculum? (b) How frequently do you use GenAI tools in course activities? Not at all importantNeverVery importantVery often52552762732402930
Figure 3. Surveyed instructors felt it was important to integrate GenAI tools into their curriculum (a), but fewer instructors reported actually doing so (b). Instructors responded on a 5-point scale. The number of responses for each option is displayed within each bar.
Stacked bar plots showing responses to two survey questions: (a) How important do you think it is to integrate GenAI tools into your curriculum?, (b) How frequently do you use GenAI tools in course activities?
@@ 4.6. Lack of Resources
0%2525%5050%7575%100100% (a.i) Do you have access to sufficient resources (funding, training, tools) to implement generative AI effectively? (a.ii) Do you have access to sufficient resources (funding, training, tools) to implement generative AI effectively? (b.i) What is your current teaching load per academic term (on average)? (MSI) (b.ii) What is your current teaching load per academic term (on average)? (non-MSI) (c) Do you feel you have the time to learn and integrate GenAI into your teaching, given your current workload? (d) How well does your institution support faculty in adopting GenAI? NoNot sureYes531122NoNot sureYes2824131-2 courses3-4 courses5-66+263810101-2 courses3-4 courses5-631288Strongly disagreeStrongly agree1534503525Very poorlyExcellent1327613623
Figure 4. A substantial number of our survey respondents expressed that they did not have the resources they needed to implement GenAI, which was especially prominent for MSI instructors (a.i, a.ii). MSI instructors were also more likely to have heavier teaching loads (b.i, b.ii). Instructors generally perceived a lack of sufficient time to learn and integrate GenAI (c), and that their institutions did not support this work (d). Instructors responded on a 5-point scale. The number of responses for each option is displayed within each bar.
Stacked bar plots showing responses to four survey questions: (a) Do you have access to sufficient resources to implement GenAI effectively?, (b) What is your current teaching load per academic term (on average)? (c) Do you feel you have the time to learn and integrate GenAI into your teaching, given your current workload?, (d) How well does your institution support faculty in adopting GenAI?
Adapting courses for GenAI required resources that many surveyed computing instructors lacked. When asked whether they had sufficient resources like funding, training, and tools to implement GenAI effectively, 53% responded No. This pattern was more pronounced among MSI instructors: 62% responded No, compared to 43% of non-MSI instructors (Figure 4a.i, a.ii). Teaching load may help explain this difference. MSI instructors were more likely to teach three or more courses per term (70% vs. 54% for non-MSIs) (Figure 4b.i, b.ii). In the most extreme cases, 10 respondents reported teaching six or more courses per academic term; all were from MSIs.
Instructors also expressed a lack of time: when asked if they had time to learn and integrate GenAI given their current workload, 62% of surveyed instructors responded Strongly Disagree, Disagree, or Neither (Figure 4c). When asked how well their institution supports faculty in adopting emerging technologies like GenAI, a nearly identical 63% of faculty responded Very Poorly, Poorly, or Neither (Figure 4d).
To understand the kinds of resources actually required to adapt a course, we examined the resources that our interviewees had when conducting emergency pedagogical design. We found that our interviewees dedicated substantial amounts of time, staff hours, and sometimes even their own money in the process of emergency pedagogical design. For example, bespoke infrastructure required ongoing development and maintenance. As mentioned in Section 4.4, P04 deployed a custom chatbot and P09 built an AI-graded explanation workflow; both of these systems required design, implementation, bug fixes, and upkeep during the term. These two instructors reported that these systems were only possible with financial support to hire software engineers and PhD students.
@@
Alternative assessments often meant a shift away from autograded work towards qualitative work, which also required more course staff to manage. As mentioned in Section 4.5, P02 reorganized his course around weekly stand-up meetings with every student. To handle this for his course of 300 students, he needed “a lot of TAs!” (P02). In practice, this course needed to recruit and train over 50 course staff members to cover approximately 300 students, a 6:1 staff to student ratio. Others emphasized the hand-grading burden for creative work, especially in courses that traditionally relied on autograder tests (P05, P08, P13).
@@
What is notable about our interviewees is that they taught at most two courses per academic term, and many had advantages such as partnerships, external funds, curriculum materials, or the ability to hire many course staff members. This pattern is striking: substantial attempts to change student-AI interaction were made by instructors with lighter teaching loads and access to staffing or funding. These accounts highlight the kinds of resources – time, staffing, funding, and infrastructure – that enabled our interviewees to make substantial course changes. Yet our survey data suggests that most computing instructors, especially those at MSIs who reported heavier teaching loads and less institutional support, do not have comparable resources. Because we view our survey sample as a more representative of the constraints faced by computing faculty, this gap indicates that many computing instructors may be unable to adopt the kinds of course adaptations described by our interviewees without additional support.
Key Insights: Instructors who successfully adapted courses for GenAI tended to have advantages like lower teaching load, staffing, and funding. Instructors who lacked these resources, especially those at MSIs, found it more difficult to effectively engage in emergency pedagogical design.